OpenStudy (anonymous):

solve: tan x = -sqrt(3). (unit circle problem)

OpenStudy (experimentx):

http://math2.org/math/trig/tables.htm

OpenStudy (anonymous):

The tangent function range is pi, not 2pi. So in one rotation of the unit circle there should be two answers. When tan(x) is negative, that means either the sine or cosine is negative. So. tan(x) = -sqrt(3) / 1 = (-sqrt(3) /2) / (1/2) x = 2pi/3 and 5pi/3 of course these answers must also include any coterminal angle.

OpenStudy (anonymous):

thanks yall!!!!

OpenStudy (anonymous):

Sure thing :)